### v1

# ARS: Automatic Routing Solver with Large Language Models

Abstract
Real-world Vehicle Routing Problems (VRPs) are characterized by a variety of practical constraints, making manual solver design both knowledge-intensive and time-consuming. Although there is increasing interest in automating the design of routing algorithms, existing research has explored only a limited array of VRP variants and fails to adequately address the complex and prevalent constraints encountered in real-world situations. To fill this gap, this paper introduces RoutBench  [^1] https://github.com/Ahalikai/ARS-Routbench, a benchmark of 1,000 VRP variants derived from 24 attributes, for evaluating the effectiveness of automatic routing solvers in addressing complex constraints. Along with RoutBench, we present the Automatic Routing Solver (ARS), which employs Large Language Model (LLM) agents to enhance a backbone algorithm framework by automatically generating constraint-aware heuristic code, based on problem descriptions and several representative constraints selected from a database. Our experiments show that ARS outperforms state-of-the-art LLM-based methods and commonly used solvers, automatically solving 91.67% of common VRPs and achieving at least a 30% improvement across all benchmarks.
Vehicle Routing Problem, Large Language Model, Automatic Heuristic Design, Benchmark.
1 Introduction
The Vehicle Routing Problem (VRP) is a fundamental Combinatorial Optimization Problem (COP) that plays a critical role in logistics, transportation, manufacturing, retail distribution, and delivery planning (Toth & Vigo, 2014; Liu et al., 2023). In these scenarios, the objective of VRPs is to efficiently allocate and plan vehicle routes to meet various requirements while minimizing overall routing costs. These requirements often include constraints such as vehicle capacities, time windows, and duration limits, resulting in numerous variants of VRPs in practical applications (Braekers et al., 2016). Existing heuristics are usually problem-specific. When the problem changes slightly (e.g., a minor modification to the requirements), a lot of effort is required for experts to redesign the heuristic to make it effective in solving the new problems (Vidal et al., 2013; Rabbouch et al., 2021; Errami et al., 2023).
Large Language Models (LLMs) have shown powerful reasoning and code-generation capabilities (Chen et al., 2021; Austin et al., 2021; Li et al., 2023). By integrating these functionalities, users can express their specific requirements in natural language, enabling the models to automatically design algorithms to address VRPs (Liu et al., 2024b). Most existing works primarily leverage the LLM to solve a small number of standard VRPs (Jiang et al., 2024; Huang et al., 2024) and cannot be applied to complex VRPs. Moreover, current methods rely on general solvers to optimize routes (Xiao et al., 2023; Zhang et al., 2024). However, solvers often struggle with standard modeling and exhibit low solving efficiency, making it difficult to handle complex constraints that arise in practical applications.
To tackle these challenges, this paper proposes a framework that uses a heuristic algorithm, developed with the assistance of LLMs, to solve the VRP variants with complex constraints. Our contributions are summarized as follows:
  • We propose ARS, a framework designed to automatically create constraint-aware heuristics based on the problem description, which enhances a backbone heuristic algorithm for route optimization, offering an adaptive framework to address the diverse routing problems expressed in natural language.
  • We introduce RoutBench, a benchmark with 1,000 VRP variants derived from 24 VRP constraints. Each variant in RoutBench is equipped with a detailed problem description, instance data, and validation code, enabling the evaluation of various routing solvers’ effectiveness in handling diverse VRP constraints.
  • We comprehensively validate our approach on commonly-used problems and RoutBench. The results show that ARS can automatically handle 91.67% of commonly-used problems, achieving at least a 30% improvement across all benchmarks compared to other LLM-based methods.
2 Problem Formulation
Vehicle Routing Problems (VRPs) involve optimizing the routes and schedules of a fleet of vehicles delivering goods or services to various locations, aiming to minimize costs while satisfying constraints like delivery windows and vehicle capacities. The VRP variants can be mathematically described as optimization problems on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where nodes $\mathcal{V}=\{0,1,\ldots,n\}$ represent depot 00 and locations $\{1,\ldots,n\}$, and edges represent the possible routes between these nodes =$\mathcal{E}=\{e_{ij},i,j\in\mathcal{V}\}$, each of them is assigned with a cost $c_{ij}$. The mathematical representation is given by:
$\min\sum_{i\in V}\sum_{j\in V}c_{ij}x_{ij},$ (1)
subject $\text{subject to}\quad\mathbf{x}\in C,$
where $\mathbf{x}=\{x_{ij}\mid i,j\in\mathcal{V},\ i\neq j\}$ represents the set of decision variables, $x_{ij}$ is a binary variable that indicates whether the route from $i$ to $j$ is used. The feasible solution space $C$ is defined by constraints. In this paper, we consider VRPs with a variety of real-world constraints such as vehicle capacity, travel distance, and time windows, thereby extending the basic VRP, as seen in the Capacitated VRP (CVRP) (Toth & Vigo, 2014) and the VRP with Time Windows (VRPTW) (Solomon, 1987). Moreover, new constraints often emerge in real-world scenarios. For example, VRP variants related to vehicle capacity include Heterogeneous VRP (HVRP), which considers vehicles with different capacities (Lai et al., 2016), Multi-Product VRP (MPVRP), which addresses the need to transport multiple types of products (Yuceer, 1997), and dynamic demands (Powell, 1986). VRP variants that incorporate these real-world constraints are more prevalent and practically significant in real-world applications
However, current methods focus on a limited range of problems and do not sufficiently address the complex and diverse constraints present in real-world scenarios. To bridge this gap and further the development of automated solutions for practical VRP variants, this paper introduces a benchmark for VRPs featuring various complex yet practical constraints. Additionally, we propose a general automatic routing solver enhanced by a large language model to effectively manage these constraints.
3 Automatic Routing Solver
Figure 1: Overview of the proposed ARS framework. The left side of the figure shows the test instance, including the problem description, corresponding data, and validation code for result verification. The right side comprises the database, generation module, and VRP solver. The generation module selects relevant constraints from the database and generates constraint-aware heuristics for the VRP solver to address different VRP variants.
Given the problem description in natural language format and the instance data for any VRP variants with one or more constraints, our proposed automatic routing solver (ARS) can automatically generate the constraint-aware heuristic to augment the backbone heuristic solver. ARS consists of three key components: 1) Pre-defined Database, 2) Constraint-aware heuristic generation, and 3) Augmented heuristic solver.
3.1 Database
We build a database, denoted as $D$, with several representative fundamental constraints to provide additional guidance for LLM-driven constraint-aware heuristic generation. Specifically, database $D$ includes a basic VRP information $(I_{0},C_{0})$ (without additional constraints) and six representative constraints $(I_{k},C_{k}),k=1,\dots,6$, each corresponds to a distinct representative constraint: Vehicle Capacity, Distance Limit, Time Windows, Pickup and Delivery, Same Vehicle, and Priority.
Each constraint example $(I_{k},C_{k})$ consists of two parts:
  • $I_{k}$: The problem description. The natural language description of the constraint.
  • $C_{k}$: The constraint feasibility checking program. It checks whether a solution belongs to the feasible solution space described by $I_{k}$. The program is given a solution and returns ’True’ if the corresponding constraint is satisfied.
3.2 Constraint-Aware Heuristic
We first select relative constraints from the database. Then we sequenticaly generate the constraint checking and scoring programs for the target VRP variants given the selected constraints and the problem description. Finally, the constraint-aware heuristic is generated based on the designed constraint checking and scoring programs.
3.2.1 Constraint Selection
Given an input problem description $I$, we instruct LLM agents to automatically select a subset of relevant constraints $S$ from the database $D$ to provide reference information for subsequent constraint program generation. There can be two cases. In the first case, LLM agents select one or more relative constraint examples. In the second case, if no constraints are recognized as related to the input $I$, the base case $(I_{0},C_{0})$, with no additional contrasts, is selected.
This step is similar to Retrieval-Augmented Generation (RAG) methods that retrieve information from external sources (Zhang et al., 2023; Zhou et al., 2022), while our candidate reference information is a pre-defined database with representative constraints.
3.2.2 Constraint Checking Program Generation
This approach works by taking a set of relevant constraints, denoted as $S$. For each constraint $C_{k}$ within $S$, we incorporate specific modifications, $\Delta C_{k}(I)$, based on the problem description $I$. This process results in the creation of new, customized constraints $C_{new}$, which are better suited to address the specific requirements of the problem. This method offers two main benefits:
  • User-Centric Design: It aligns with how users typically work by enhancing existing constraints. This allows users to refine their specific requirements without the need to develop a complete problem definition from the beginning.
  • Efficient Processing by LLMs: It helps LLM agents focus on these new, tailored constraints. This reduces unnecessary complexity and enhances the relevance and accuracy of the solutions provided by the models.
3.2.3 Constraint Scoring Program Generation
In practice, heuristic algorithms operate within a variable space that is typically much larger than the feasible solution space. This discrepancy presents significant challenges in identifying feasible solutions. To address this, heuristic methods may permit the presence of infeasible solutions during the search process, as this allows for the evaluation of potential improvements in solution quality (Deb, 2000; Máximo & Nascimento, 2021).
Therefore, to effectively integrate $C_{\text{new}}$ into the solution process, we utilize LLM agents to generate a violation score function guided by the constraint checking program. This score function quantifies the degree of constraint violation, thereby establishing a method for handling constraints and identifying high-quality solutions. It aids in systematically assessing and managing constraint violations, facilitating the search for feasible and high-quality solutions within the expansive variable space.
3.2.4 Constraint-Aware Heuristic
We present the Constraint-Aware Heuristic (CAH) based on the constraint checking (Checker) and scoring programs (Scorer). As seen in Algorithm 1, the constraint handling method evaluates whether a new solution ${s}_{new}$ improves upon an old solution ${s}_{old}$. It first verifies whether ${s}_{new}$ is feasible (line 1). If ${s}_{new}$ is feasible and ${s}_{old}$ is infeasible (line 2), or both are feasible but ${s}_{new}$ results in a smaller travel distance (line 4), ${s}_{new}$ is accepted. If both are infeasible, the algorithm compares the violation scores of both solutions (line 8) and selects the one with the lower score. This approach allows infeasible solutions to evolve gradually toward feasibility while minimizing the overall travel distance.
Algorithm 1 Constraint-Aware Heuristic
0:  New solution ${s}_{new}$, Old solution ${s}_{old}$
1:  if Checker(${s}_{new}$) is feasible then
2:     if Checker(${s}_{old}$) is infeasible then
3:        return true;
4:     else if Cost(${s}_{new}$) $<$ Cost(${s}_{old}$then
5:        return true;
6:     end if
7:  else if Checker(${s}_{old}$) is infeasible then
8:     if Scorer(${s}_{new}$) $<$ Scorer(${s}_{old}$then
9:        return true;
10:     end if
11:  end if
3.3 Augmented Heuristic Solver
The solver has a general single-point-based search backbone heuristic framework. The automatically generated constraint handling codes are used for this backbone heuristic framework.
3.3.1 Backbone Heuristic Framework
The backbone heuristic solver mainly consists of 1) destroy&repair and 2) local search.
In the destroy phase, we employ multiple operators, including random removal and string removal (Christiaens & Vanden Berghe, 2020), to selectively remove customers or partial routes from the current solution. The choice of destroy operators is determined using a roulette wheel selection mechanism, which assigns higher probabilities to operators that performed well in previous iterations. The repair phase reinserts removed customers into the solution using a greedy repair operator, aiming to construct a feasible solution with shorter routes. Specific details of these operators are provided in the Appendix B.
Following the destroy and repair phases, the solution undergoes a local search process to further refine its quality. We utilize a set of local search operators, including 2-OPT, SWAP, and SHIFT, and the best solution found among these operators is selected. To avoid premature convergence, the Record-to-Record Travel (RRT) criterion is applied (Dueck, 1993; Santini et al., 2018), allowing the acceptance of slightly worse solutions within a predefined threshold, thus maintaining a balance between intensification and diversification.
The entire process is iteratively repeated until a termination condition is met, such as a time limit.
3.3.2 Constraint-Aware Heuristic
Throughout the search, the constraint-aware heuristic plays a critical role in evaluating and guiding solutions. It is used in the recreate step and all local search operators. By effectively navigating infeasible regions and gradually improving solution feasibility, it ensures that the solver not only satisfies user requirements but also effectively minimizes the total route length.
Table 1: A classification of common VRP variants is presented based on six constraint types, with abbreviations provided for both the constraints and VRP variants in parentheses. Each constraint type includes four distinct variants, with examples provided for each, resulting in a total of 24 constraints. All of these constraints have already attracted attention from researchers.
Basic Constraints VRP Variants
Vehicle Capacity (C) Capacitated VRP (CVRP) (Vidal, 2022; Luo et al., 2023)
Heterogeneous CVRP (HCVRP) (Lai et al., 2016)
Multi-Product VRP (MVRP) (Yuceer, 1997)
Dynamic CVRP (DCVRP) (Powell, 1986)
Distance Limit (L) VRP with Distance Limit (VRPL) (Laporte et al., 1985; Li et al., 1992)
Heterogeneous VRPL (HVRPL) (Lee et al., 2021)
Recharging VRP (RVRP) (Conrad & Figliozzi, 2011; Erdoğan & Miller-Hooks, 2012)
Dynamic VRPL (DVRPL) (Suzuki, 2011; Khouadjia et al., 2012; Qian & Eglese, 2016)
Time Windows (TW) VRP with time windows (VRPTW) (Solomon, 1987)
Heterogeneous VRPTW (HVRPTW) (Ren et al., 2010; Vidal et al., 2014)
VRP with Multiple Time Windows (VRPMTW) (Belhaiza et al., 2014)
Dynamic VRPTW (DVRPTW) (Ghiani et al., 2003; Pillac et al., 2013)
Pickup and Delivery (PD) VRP with mixed pickup and delivery (VRPMPD) (Avci & Topaloglu, 2015)
Heterogeneous VRPMPD (HVRPMPD) (Avci & Topaloglu, 2016)
Multi-Product VRPMPD (MVRPMPD) (Zhang et al., 2019)
Dynamic VRPMPD (DVRPMPD) (Gendreau et al., 2006)
Same Vehicle (S) VRP with Same Vehicle Constraint (VRPSVC) (Kumar & Panneerselvam, 2012)
Clustered VRP (CluVRP) (Battarra et al., 2014; Islam et al., 2021)
VRP with Sequential Ordering (VRPSO) (Escudero, 1988; Gambardella & Dorigo, 2000)
VRP with Incompatible Loading Constraint (VRPILC) (Wang et al., 2015)
Priority (P) Precedence constrained VRP (PVRP) (Kubo & Kasugai, 1991)
VRP with Relaxed Priority Rules (VRPRP) (Doan et al., 2021)
VRP with Multiple Priorities (VRPMP) (Yang et al., 2015)
VRP with d-Relaxed Priority Rule (VRP-dRP) (Dasari & Singh, 2023; dos Santos Teixeira & de Araujo, 2024)
4 RoutBench
Recent advancements in LLMs have opened up new possibilities for automatically generating routing solvers to address different VRP variants (Xiao et al., 2023). These methods have shown promising automation capabilities and have been applied to a few VRP variants (Chen et al., 2023). However, their evaluations are often constrained to some simple problems, leaving a significant gap in assessing their generalization ability. There is yet to be a VRP benchmark that can evaluate the generalization ability of these LLM-based methods, especially under complex yet common constraints in real-world scenarios.
Thus, we propose RoutBench, a benchmark dataset that includes 1,000 VRP variants derived from 24 constraints, each has three different sizes: 25, 50, and 100. As shown in Table 1, 4 variants are selected for each of the six constraint types: Vehicle Capacity (C), Distance Limits (L), Time Windows (TW), Pickup and Delivery (PD), Same Vehicle (S) and Priority (P). These constraint types are representative due to their practical significance and theoretical challenges, which already have attracted researchers’ attention. These 1,000 problem instances can be divided into simple and hard based on the number of constraints.
  • RoutBench-S: 500 simple VRP variants with 3 or fewer constraints
  • RoutBench-H: 500 hard VRP variants with 4 or more constraints
Each problem instance is comprised of three components: 1) the problem description, which is a natural language explanation of the problem; 2) the instance data, including the geometric positions of nodes and the constraint parameters; and 3) the validation code, used to confirm whether a solution adheres to user requirements and satisfies all constraints.
By offering a comprehensive and standard dataset for VRP variants with diverse and practical constraints, RoutBench provides a valuable test benchmark for advancing research on vehicle routing algorithms, fostering the development of practical routing solvers.
4.1 Dataset Construction
The RoutBench is constructed by combinations of six basic constraint types: vehicle capacity, distance limits, time windows, pickup and delivery, same vehicle constraints, and node priority. We design four representative variants derived from each basic constraint type that incorporate variations such as heterogeneous vehicle fleets, multidimensional resource limits, dynamic changes, and others. To produce one problem instance, we first pick one basic constraint combination and then pick one real constrain variant for each chosen basic constraint type. Notice that Vehicle Capacity (C) and Pickup and Delivery (PD) are mutually exclusive, since PD is essentially a derivative of C. The total number of problem combinations is :
$N_{\text{combinations}}=\sum_{k=1}^{6}\binom{6}{k}\cdot 4^{k},$ (2)
where $\binom{6}{k}$ represents the number of ways to select $k$ constraints from the six types, and $4^{k}$ accounts for the four variations per constraint. After excluding combinations where vehicle capacity and pickup and delivery coexist, the total number of feasible combinations is reduced to 5624. From these, 1000 unique problem instances are uniformly sampled based on the order of all combinations, ensuring even coverage across the solution space. To balance complexity, 500 problems are sampled with three or fewer constraints, while the remaining 500 include more than three constraints.
Each problem instance includes detailed parameters such as customer locations, demands, time windows, and other relevant attributes, with data generated using the Solomon C103 dataset as a base. Problem sizes include 25, 50, and 100 nodes. To ensure every problem has at least one feasible solution, the VRP solver in this study directly incorporates a demand verification code into its Selection Strategy. This strategy guides the search process during problem generation, ensuring that each problem is constructed with a valid and feasible solution.
The detailed descriptions of the 24 problem instances are provided in Appendix D. These examples illustrate the diversity of problem settings and serve as a representative subset of the dataset’s broader scope. By systematically combining constraints, leveraging validation mechanisms, and ensuring feasibility, RoutBench offers a diverse and reliable dataset for benchmarking VRP solvers.
4.2 Analysis
This section analyzes the distribution of problem types and complexities in the RoutBench dataset, focusing on the 48-problem subset and RoutBench, and their relationship to the full set of 5624 feasible problems.
The distribution of problems by the number of constraints is shown in Figure 2. Among the 48 problems, one is a simple VRP without any constraints, while the remaining 47 include one to five constraints. In RoutBench, the distribution reflects the proportions of the full set of 5624 problems. Specifically, problems with three constraints are the most common in the dataset, as three-constraint problems dominate the total number of problems with three or fewer constraints. For more complex problems, those with four and five constraints appear in similar proportions, ensuring a balanced representation of high-complexity scenarios.
Figure 3 shows the distribution of problem types across RoutBench. Most problem types, such as VRPTW, HVRPL, and CluVRP, are well-represented. However, problems involving vehicle capacity constraints (e.g., CVRP, HCVRP) and pickup and delivery operations (e.g., VRPMPD, HVRPMPD) are less frequent. This is because these two categories are mutually exclusive, and the two types do not coexist in the dataset.
Overall, the RoutBench dataset achieves a diverse and balanced representation of problem types and complexities. The 48-problem subset provides a concise overview of simpler cases, while RoutBench captures a wide range of scenarios.
Figure 2: Analysis of problem distribution across common problems, RoutBench, and all problems. Among the common problems, one is a simple VRP without any constraints.
Table 2: The performance comparison between ARS and seven LLM-based methods on common problems and RoutBench. The best results among these methods are highlighted in grey.
Methods Common Problems RoutBench
RoutBench-S RoutBench-H
SR $\uparrow$ RER $\downarrow$ SR $\uparrow$ RER $\downarrow$ SR $\uparrow$ RER $\downarrow$
Standard Prompting 41.67% 6.25% 37.60% 8.20% 11.60% 15.80%
CoT 37.50% 8.33% 37.40% 9.60% 13.40% 14.80%
Reflection 45.83% 6.25% 41.80% 4.60% 15.20% 7.60%
PHP 33.33% 10.42% 32.60% 12.00% 11.20% 17.40%
CoE 37.50% 2.08% 34.20% 8.00% 11.40% 12.20%
Self-debug 43.75% 0.00% 34.60% 4.00% 10.80% 7.60%
Self-verification 43.75% 2.08% 34.00% 5.20% 15.60% 9.20%
ARS (Ours) 91.67% 0.00% 73.20% 5.20% 46.80% 11.80%
5 Experiments
The experiments are conducted on common benchmark problems and RoutBench, which include 48 and 1,000 VRP variants, respectively. These variants cover a wide range of real-world problem types, such as heterogeneous, dynamic, multi-commodity, and others.
To evaluate the performance of these LLM-based methods in handling different VRP variants, we assess their ability to generate correct programs within our solver. To analyze the impact of different solvers on the generated programs, we compare our solver with other solvers (e.g., Gurobi and OR-Tools) to investigate their influence. Additionally, we included experiments with other open-source and proprietary LLMs (e.g., DeepSeek v3 and LLaMA 3.1 70B) to explore their impact on program development.
Figure 3: The frequency of constraint usage is analyzed for both RoutBench-S and RoutBench-H.
Table 3: Comparison of program development performance across different solvers for common problems using standard prompting. The best results among these solvers are highlighted in grey.
Solvers Common Problems
SR $\uparrow$ RER $\downarrow$
Standard Prompting (OR-tools) 12.50% 58.33%
Standard Prompting (Gurobi) 14.58% 41.67%
Standard Prompting (Ours) 41.67% 6.25%
5.1 Comparison with LLM-Based Methods
The first experiment aims to evaluate the ability of ARS to produce successful programs. We compare our proposed ARS with seven LLM-based methods, including Standard Prompting, Chain of Thought (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), Progressive-Hint Prompting (PHP) (Zheng et al., 2023), Chain-of-Experts (CoE) (Xiao et al., 2023), Self-debug (Chen et al., 2023), and Self-verification (Huang et al., 2024). To focus on program development, all methods are implemented using our backbone algorithm framework to handle various VRP variants.
This experiment is conducted on the common problems and RoutBench. We use two metrics to evaluate the results. The Success Rate (SR) measures the proportion of the program where generated solutions pass the validation process. The Runtime Error Rate (RER) quantifies the percentage of programs that fail due to issues such as runtime errors, incorrect API usage, or syntax mistakes.
As shown in Table 2, ARS significantly outperforms other LLM-based methods in generating correct programs for both common problems and RoutBench. It is evident that these compared methods exhibit unsatisfactory performance across all problems, particularly on the RoutBench-H problems, where their SR is merely around 10%. None of the compared algorithms achieved an SR of 50% or higher. In contrast, ARS achieved an SR of 91.67% on the common problems, successfully addressing 44 out of 48 VRPs. Moreover, on RoutBench, ARS outperformed all seven other LLM-based methods by at least 30% in terms of SR. These results highlight ARS’s competitiveness in addressing complex VRP variants.
5.2 Comparison with Different Solvers
In the previous experiment, to focus on generating correct programs, all methods use the same solving framework to handle VRP variants. To explore the impact of different solving tools on program development, this experiment uses Standard Prompting with three solving tools (OR-tools, Gurobi, and our backbone) for common problems.
As shown in Table 3, both SR and RER achieve significantly better results when using our framework compared to OR-tools and Gurobi. Under standard prompting, our framework achieves an SR of 41.67%, whereas OR-tools and Gurobi reach only about 10% SR. This is because, in our framework, the LLM only needs to complete the constraint codes in Python, whereas the use of solvers is restricted by their specific implementations and requires highly standardized modeling. For the same VRP variants, our solving framework simplifies program development considerably compared to these solvers. Examples of program codes are provided in the appendix F.
To evaluate the performance of our solver on different VRP variants, we also conducted a comparative study with two other solvers across three common VRP variants (e.g., CVRP, CVRPTW, and CVRPL). The experiments are conducted on instances with 25 and 50 nodes, with time limits of 50 seconds and 5 minutes, respectively.
As can be seen from Table 4, our solver achieved the optimal results in many instances. Notably, it finds better solutions than the other two solvers in CVRP and CVRPL instances. These findings highlight the significant potential of heuristic algorithms in solving VRP variants effectively.
Table 4: Performance analysis of different solvers on three common VRP variants. The table presents the gaps compared to the results obtained by Gurobi. The best results among these solvers are highlighted in grey.
CVRP
Test sets N=25, C=200 N=50, C=200
Solvers Obj. $\downarrow$ Gap Obj. $\downarrow$ Gap
OR-tools 223.6 18.24% 420.4 16.33%
Gurobi 189.1 0.00% 361.4 0.00%
Ours 186.9 -0.94% 358 -1.16%
CVRPTW
Test sets N=25, C=200 N=50, C=200
Solvers Obj. $\downarrow$ Gap Obj. $\downarrow$ Gap
OR-tools 192.2 1.00% 392.0 8.47%
Gurobi 190.3 0.00% 361.4 0.00%
Ours 190.3 0.00% 361.4 0.00%
CVRPL
Test sets N=25, C=200, L=120 N=50, C=200, L=120
Solvers Obj. $\downarrow$ Gap Obj. $\downarrow$ Gap
OR-tools 221.3 17.65% 393.4 9.04%
Gurobi 188.1 0.00% 360.8 0.00%
Ours 186.9 -0.64% 358 -0.78%
(a) RoutBench-S
(b) RoutBench-H
Figure 4: The performance of standard prompting and ARS is compared across various LLMs. ARS demonstrates compatibility with different models and shows clear advantages in RoutBench.
5.3 Evaluation with Different LLMs
We further investigate the impact of using different LLMs for generating correct programs to solve VRP variants under standard prompting and ARS. The LLMs evaluated include GPT-3.5-Turbo, GPT-4o, DeepSeek-V3, and LLaMA-3.1-70B.
As shown in Figure 4, the success rate of solving VRP variants varies across different LLMs. All methods benefit from more advanced LLMs, leading to improved accuracy. The results indicate that DeepSeek-V3 is the most effective LLM for handling VRP variants among these LLMs. When using DeepSeek-V3, ARS achieves an SR of 77.20% on simple problems in RoutBench. The improvement is even more pronounced for RoutBench-H problems, where ARS attains an SR of 61.60%.
5.4 Ablation Study
An ablation study is conducted to evaluate the program development component of ARS. To analyze the impact of our database and constraint selection step on ARS, experiments are performed by removing each of these components separately.
As shown in Table 5, removing the constraint selection step leads to a decrease in the SR of ARS. This is because, without constraint selection, ARS acquires all six representative constraints, which are not necessarily all relevant to the current VRP. Consequently, this irrelevant information may mislead the LLM.
Table 5: Ablation study on the effect of component removal on ARS’s ability to generate correct programs. The best results among these methods are highlighted in grey.
Methods Common Problems
SR $\uparrow$ RER $\downarrow$
ARS (Full) 91.67% 0.00%
w/o Constraint Selection 62.50% 2.08%
w/o Database 35.42% 12.50%
Compared to constraint selection, removing the database has a more significant impact on the performance of ARS in generating correct programs. As previously discussed, ARS can be refined based on relevant constraints, thereby focusing on the distinctive aspects of a VRP variant. However, without reference to relevant constraints, the LLM must independently generate all the constraints, imposing a higher requirement on the LLM.
The results of this ablation study demonstrate that ARS achieves optimal program development performance only when both components are utilized.
6 Conclusion
In this paper, we have proposed ARS, a framework that leverages LLMs to automatically design constraint-aware heuristics to enhance a heuristic routing optimization framework for tackling VRP variants with complex and practical constraints. Moreover, we have developed and presented RoutBench, a comprehensive benchmark consisting of 1,000 VRP variants derived from 24 VRP constraints. Each variant includes a problem description, instance data, and validation code, enabling standard evaluation of routing solvers. The experimental evaluations of ARS have been particularly promising, demonstrating a superior performance over existing LLM-based methods and commonly used solvers. Specifically, ARS achieved a remarkable success rate of 91.7% in automatically solving common VRP instances and registered at least a 30% improvement across all tested benchmarks. In the future, we would like to extend the ARS framework to address other combinatorial optimization challenges, including scheduling and bin packing.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related works
A.1 Heuristics for VRPs
The Vehicle Routing Problem (VRP) is a classical combinatorial optimization problem that seeks optimal routes for vehicles to serve customers under constraints like capacity and time windows (Dantzig & Ramser, 1959). Variants such as the Capacitated VRP (CVRP) (Lysgaard et al., 2004), VRP with Time Windows (VRPTW) (Solomon, 1987), and Multi-Depot VRP (MDVRP) (Yuceer, 1997) have been extensively studied.
To address VRP’s computational challenges, heuristic methods are widely used. Local Search (LS) iteratively improves solutions using operators like 2-opt or 3-opt (Lin, 1965), while Iterated Local Search (ILS) introduces perturbations to escape local optima (Lourenço et al., 2003). Simulated Annealing (SA) explores worse solutions probabilistically to avoid stagnation (Kirkpatrick et al., 1983), and Tabu Search (TS) prevents revisiting solutions using a tabu list (Glover, 1989). Variable Neighborhood Search (VNS) systematically adjusts neighborhood structures for better exploration (Mladenović & Hansen, 1997).
Adaptive Large Neighborhood Search (ALNS) selects neighborhood operators adaptively based on performance (Ropke & Pisinger, 2006). Recent ALNS variants integrate machine learning to predict effective operators (Mara et al., 2022) and reinforcement learning to optimize selection policies (Wang et al., 2024). Hybrid ALNS approaches, such as combining ALNS with branch-and-price, have further improved performance on constrained VRPs (Vidal, 2022).
In summary, these heuristic advancements have significantly promoted VRP research by enabling efficient solutions to increasingly complex and large-scale problem instances.
A.2 LLMs for VRPs
Recent advancements in large language models (LLMs) have opened new possibilities for solving vehicle routing problems (VRPs) by leveraging their ability to process and embed complex tasks through natural language understanding (Huang et al., 2024). These models have been applied to classical optimization problems such as the traveling salesman problem (TSP) and CVRP. In particular, LLMs, combined with evolutionary computation (EC), have demonstrated potential in automating the discovery of heuristics for VRPs (Liu et al., 2024a). For instance, guided local search and other algorithmic frameworks can be evolved with the assistance of LLMs, albeit requiring significant domain-specific expertise and computational resources for iterative model invocation (Wang et al., 2024).
Another line of research emphasizes the transformation of textual problem descriptions into mathematical formulations and executable code that can be processed by external solvers (Tang et al., 2024). This approach benefits from LLMs’ ability to interpret user queries and generate structured outputs, enabling the automation of optimization tasks. Furthermore, multi-agent systems have been introduced to coordinate LLM-based agents for tasks such as problem formulation, programming, and evaluation (Xiao et al., 2023). While promising, these methods are inherently limited by the scope of the knowledge encoded within the pre-trained models, which restricts their applicability to novel or highly complex problems (Zhang et al., 2024).
To address these challenges, emerging studies focus on directly generating executable programs for VRPs by integrating LLMs with external knowledge sources. This approach bypasses the intermediate step of mathematical model formulation and instead leverages both the intrinsic reasoning capabilities of LLMs and external references to enhance problem-solving efficacy (Huang et al., 2024). Such integration aims to reduce reliance on domain-specific expertise while expanding the range of solvable problems, marking a shift towards more generalized and adaptive optimization frameworks (Romera-Paredes et al., 2024).
Appendix B Operator
Local search operators are critical components of heuristic methods, designed to explore the neighborhood of a solution and improve its quality. The commonly used operators are as follows:
2-opt Operator. The 2-opt operator is a classical approach originally developed for the Traveling Salesman Problem (TSP). It works by removing two non-adjacent edges in the solution and reconnecting them in a different way, thereby altering the order of nodes. If the new configuration reduces the total cost, it is accepted as an improved solution.
Swap Operator. The Swap operator exchanges the positions of two elements in the solution, generating a new configuration. This simple yet effective operator promotes diversity in the solution space and helps avoid local optima.
Shift Operator. The Shift operator moves an element from its current position to another position within the same solution or a different one. This adjustment redistributes the structure of the solution, enabling further exploration of the search space and improving solution quality.
Destroy Operator. The Destroy operator partially disrupts the current solution by selectively removing some elements, breaking its local optimality and enabling exploration of new regions in the search space. Two common implementations are Random Removal and String Removal:
  • Random Removal: This method removes elements uniformly at random, without relying on specific heuristics, making it a simple yet effective way to diversify the search process.
  • String Removal: This method targets sequences of consecutive or related elements (strings), such as partial routes or groups of customers (Christiaens & Vanden Berghe, 2020). It begins by randomly selecting a ”center” customer and removing a string of nearby customers from the route. The string size is randomly determined, constrained by the average route size and a predefined maximum. If constraints on the number of disrupted routes or previously disrupted routes are met, further removal is skipped.
Repair Operator. The Repair operator complements the Destroy operator by reinserting removed elements to reconstruct a complete solution, guided by optimization objectives. A commonly used implementation is the Greedy Repair:
  • Greedy Repair: This method reinserts removed elements one by one, selecting at each step the position that minimizes the objective function. By prioritizing optimal insertions, it balances solution quality with search efficiency.
These operators, through targeted adjustments, provide the foundation for heuristic and metaheuristic algorithms, enabling efficient exploration and exploitation of the solution space.
Appendix C Prompts of ARS
Automatic Routing Solver (ARS) is designed to perform three key steps leveraging the capabilities of LLMs: Constraint Selection, Constraint Checking Program Generation, and Constraint Scoring Program Generation. In this section, we describe the prompt engineering involved in each step. These prompts are constructed based on user inputs and several representative constraints stored in the database to generate the Constraint Checking Program and the Constraint Scoring Program. Variable information, such as user inputs and constraint descriptions, is highlighted in blue for clarity.
Step 1: Constraint Selection. In the first step, the ARS identifies constraints relevant to the user’s input from the database. This step processes the user input and matches it against the constraints stored in the database. If relevant constraints are found, they are selected for further processing. Otherwise, LLM outputs “No Relevant Constraint”. This step ensures that only the constraints pertinent to the user’s query are considered in subsequent steps.
Prompt for Constraint Selection For the description in the VRP problem, identify and provide the relevant constraint types from the following list: {constraint_description} According to the user input: {user_input} If no constraint types match the user input, respond with: ”No Relevant Constraint”. Do not give additional explanations.
Step 2: Constraint Checking Program Generation. Based on the relevant constraints selected in the previous step, the ARS uses the LLM to generate new constraints by taking the selected constraints as references. Specifically, the selected constraints are provided as input to the LLM, which then generates new constraints tailored to the user’s input.
Prompt for Constraint Checking Program Generation As a Python algorithm expert, please implement a function to check the constraints for the vehicle routing problem (VRP) based on the provided description and relevant code. User input: {user_input} Relevant Examples: {related_constraints_and_codes} Do not give additional explanations.
Step 3: Constraint Scoring Program Generation. In the final step, the ARS generates a Constraint Scoring Program based on the Constraint Checking Program created in the previous step. This scoring program evaluates the degree to which the constraints are satisfied, providing a quantitative measure of constraint adherence.
Prompt for Constraint Scoring Program Generation As a Python algorithm expert, please implement a function to calculate the constraint violation score for the Vehicle Routing Problem (VRP) based on the given constraints. Function Template: {function_template} Constraints Description: {constraints_description} Check Constraints Code: {related_constraints_and_codes} Do not give additional explanations.
Appendix D VRP Variants
The common VRP variants may involve several representative constraints, such as vehicle capacity, distance limits, time windows, pickup and delivery, same vehicle, and priority. As shown in Table 6, these six representative constraints can be combined to generate 48 common problems, with the exception of cases where vehicle capacity and pickup and delivery appear together.
Table 6: The 48 common problems constructed by six representative constraints.
Vehicle Distance Time Pickup and Same Priority
Capacity Limit Windows Delivery Vehicle
VRP
PVRP $\surd$
VRPS $\surd$
PVRPS $\surd$ $\surd$
VRPPD $\surd$
PVRPPD $\surd$ $\surd$
VRPPDS $\surd$ $\surd$
PVRPPDS $\surd$ $\surd$ $\surd$
VRPTW $\surd$
PVRPTW $\surd$ $\surd$
VRPSTW $\surd$ $\surd$
PVRPSTW $\surd$ $\surd$ $\surd$
VRPPDTW $\surd$ $\surd$
PVRPPDTW $\surd$ $\surd$ $\surd$
VRPPDSTW $\surd$ $\surd$ $\surd$
PVRPPDSTW $\surd$ $\surd$ $\surd$ $\surd$
VRPL $\surd$
PVRPL $\surd$ $\surd$
VRPLS $\surd$ $\surd$
PVRPLS $\surd$ $\surd$ $\surd$
VRPPDL $\surd$ $\surd$
PVRPPDL $\surd$ $\surd$ $\surd$
VRPPDLS $\surd$ $\surd$ $\surd$
PVRPPDLS $\surd$ $\surd$ $\surd$ $\surd$
VRPLTW $\surd$ $\surd$
PVRPLTW $\surd$ $\surd$ $\surd$
VRPLSTW $\surd$ $\surd$ $\surd$
PVRPLSTW $\surd$ $\surd$ $\surd$ $\surd$
VRPPDLTW $\surd$ $\surd$ $\surd$
PVRPPDLTW $\surd$ $\surd$ $\surd$ $\surd$
VRPPDLSTW $\surd$ $\surd$ $\surd$ $\surd$
PVRPPDLSTW $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
CVRP $\surd$
PCVRP $\surd$ $\surd$
CVRPS $\surd$ $\surd$
PCVRPS $\surd$ $\surd$ $\surd$
CVRPTW $\surd$ $\surd$
PCVRPTW $\surd$ $\surd$ $\surd$
CVRPSTW $\surd$ $\surd$ $\surd$
PCVRPSTW $\surd$ $\surd$ $\surd$ $\surd$
CVRPL $\surd$ $\surd$
PCVRPL $\surd$ $\surd$ $\surd$
CVRPLS $\surd$ $\surd$ $\surd$
PCVRPLS $\surd$ $\surd$ $\surd$ $\surd$
CVRPLTW $\surd$ $\surd$ $\surd$
PCVRPLTW $\surd$ $\surd$ $\surd$ $\surd$
CVRPLSTW $\surd$ $\surd$ $\surd$ $\surd$
PCVRPLSTW $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Table 7: Examples of constraint descriptions for 24 VRP variants.
Problem Problem Example
CVRP The total load of each route must not exceed the vehicle capacity.
HCVRP The total load of each route must not exceed the vehicle capacity. Specifically, there should be at least 3 routes where the total load is less than 100 units.
MCVRP The total load of each route must not exceed the vehicle capacity. Additionally, nodes [12, 14] require deliveries of [70, 80] units of a new type of goods. The maximum load capacity for this type of goods on each route is 100 units, and problem data excludes information about new goods.
DCVRP The total load of each route must not exceed the vehicle capacity. Specifically, for node [19], its base demand is augmented by 5 times the square root of the accumulated travel distance from the depot [0] to that node.
VRPL Each route must not exceed 150 units in length.
HVRPL Each route must not exceed 200 units in length, and at least three routes must have a total length of less than 150 units.
RVRP After visiting node [17], the vehicle’s remaining allowable travel distance for that route is reset to 150 units. At each node, the remaining driving distance cannot be negative.
DVRPL Each route must not exceed 200 units in length. The vehicle’s remaining range decreases with each visit. After visiting node [17], the remaining range will be halved.
VRPTW The arrival time at each node must meet its specified time window.
HVRPTW The arrival time at each node must meet its specified time window. Specifically, one route must have its start time is 300, while all other routes start with time 0.
VRPMTW The arrival time at each node must meet its specified time window. For node [4], in addition to its original time window, an additional time window of [900, 950] is also available.
DVRPTW The arrival time at each node must meet its specified time window. For node [18], the service time dynamically increases by the amount of time from the start of its time window to the arrival time.
VRPMPD At Node [24], 10 units of goods are picked up. Unlike delivery, visiting a pickup node reduces the vehicle’s available capacity.
HVRPMPD At Node [24], 10 units of goods are picked up. Unlike delivery, visiting a pickup node reduces the vehicle’s available capacity. Specifically, there should be at least 3 routes where the total load is less than 100 units.
MVRPMPD Nodes [12, 14] require deliveries of [70, 80] units of a new type of goods. At Node [24], I pick up 20 units of these goods and 10 units of original goods. Before pick up, it needs to check whether sufficient goods have been delivered. Both types of goods are stored separately, with a maximum load of 100 units for new goods on each route, and problem data excludes information about new goods.
DVRPMPD At Node [24], 10 units of goods are picked up, along with an additional amount calculated as 5 times the square root of the accumulated travel distance from the depot [0] to this node. Unlike delivery, visiting a pickup node reduces the vehicle’s available capacity.
VRPSVC Nodes [13, 23] must be on the same route.
CluVRP Nodes [7, 10] must be on the same route, and these nodes must be visited consecutively.
VRPSO Nodes [13, 23] must be on the same route, and node [13] must be visited before node [23].
VRPILC Nodes [7, 8] must not be on the same route.
PVRP Nodes [5, 7] are priority points.
VRPRP Node [8] is the priority node and must be one of the first three positions in at least one route.
VRPMP Nodes [7, 5, 3] are priority nodes with strictly decreasing priority levels: [7] (highest), [5], and [3] (lowest). Higher-priority nodes must be visited before lower-priority ones and other nodes.
VRP-dRP Nodes [7, 5, 3] follow the d-relaxed priority rule with decreasing priority: [7] (highest), [5], and [3] (lowest). Each node can be visited within its level or one level later, but no lower-priority node can be visited more than one level early. Other nodes are non-priority.
Appendix E Experimental Details
Experiments are performed on a computer with an Intel Xeon Gold 6248R Processor (3.00 GHz), 128 GB system memory, and Windows 10.
In RoutBench, the ARS heatmaps illustrate the frequency of simultaneous errors encountered when solving composite Vehicle Routing Problems (VRPs). The horizontal and vertical axes correspond to 24 specific VRPs, with each cell representing the total number of errors occurring when solving a composite problem that includes both the row and column problems. The diagonal values indicate the total number of errors for individual problems, reflecting their inherent difficulty.
Figure 5: The heatmap of ARS within RoutBench-S shows the number of times errors occur simultaneously for the corresponding row (horizontal axis) and column (vertical axis). The diagonal values represent the number of errors for each individual problem.
In the RoutBench-S, the number of errors in priority problems is significantly higher than in other types of problems. This may be due to the inability of LLMs to adequately understand and handle priority issues. When transitioning to the RoutBench-H, the four time-window-related problems exhibit a significantly higher number of errors compared to other problem types. This suggests that time-window problems are inherently more complex. In contrast, regardless of whether the problems are RoutBench-S or RoutBench-H, our algorithm performs exceptionally well in terms of modeling success rates for capacity constraints and return point constraints.
Figure 6: The heatmap of ARS within RoutBench-H shows the number of times errors occur simultaneously for the corresponding row (horizontal axis) and column (vertical axis). The diagonal values represent the number of errors for each individual problem.
Appendix F Examples of Program Codes
To better analyze different methods for solving VRPs, we provide code examples for three approaches: our solver, Gurobi, and OR-Tools. As a case study, these methods are applied to the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) to illustrate their respective requirements and complexities.
From the provided code examples, it is clear that our solver requires the least amount of code to be generated by an LLM. This is attributed to the simplicity and flexibility of our method, which avoids the need for verbose or overly rigid programming constructs. In contrast, the code examples for Gurobi and OR-Tools are more extensive, primarily due to their reliance on strict syntax rules and detailed configurations.
Our Solver Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Args:
4 solution (VrpState): An object representing the VRP solution.
5 - problem_data (dict): A dictionary with problem details:
6 - "edge_weight": A 2D NumPy array of distances between nodes.
7 - "demand": A 1D NumPy array of node demands (0 for the depot).
8 - "capacity": Maximum load capacity per vehicle.
9 - "service_time": A 1D NumPy array of node service times (0 for the depot).
10 - "time_window": A list of [earliest start, latest end] time windows for servicing each node.
11 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
12
13 Returns:
14 bool: True if the solution satisfies the constraints, False otherwise.
15
16 Note:
17 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
18 """
19 for route in solution.routes:
20 # Check Vehicle Capacity Constraint
21 total_demand = sum(solution.problem_data["demand"][node] for node in route)
22 if total_demand > solution.problem_data["capacity"]:
23 return False
24
25 # Check Time Windows Constraint
26 current_time = 0 # Start at time 0
27 tour = [0] + route + [0] # Add depot at the beginning and end of the route
28
29 for idx in range(1, len(tour)):
30 arrive_time = current_time + solution.problem_data[’edge_weight’][tour[idx - 1]][tour[idx]]
31 wait_time = max(0, solution.problem_data[’time_window’][tour[idx]][0] - arrive_time) # Wait if early
32 current_time = arrive_time + wait_time
33
34 tw_start, tw_end = solution.problem_data[’time_window’][tour[idx]]
35 if current_time > tw_end:
36 return False
37
38 # Add the service time for the current node after arriving and waiting
39 current_time += solution.problem_data[’service_time’][tour[idx]]
40
41 return True
Gurobi Code:
1from gurobipy import Model, GRB, quicksum
2from typing import List
3
4def find_feasible_routes(solution: VrpState) -> List[List[int]]:
5 """
6 Finds feasible routes for a CVRP problem using an optimization solver.
7
8 Args:
9 solution (VrpState): An object representing the VRP solution.
10 - problem_data (dict): A dictionary with problem details:
11 - "demand": A 1D NumPy array of node demands (0 for the depot).
12 - "vehicles": Maximum vehicle.
13 - "capacity": Maximum load capacity per vehicle.
14 - "edge_weight": A 2D NumPy array or matrix of distances between nodes.
15
16 Returns:
17 routes (list): A list of optimized routes, where each route is a list of node IDs (integers, excluding the depot node 0).
18 """
19 data = solution.problem_data
20 num_nodes = len(data["demand"])
21 demand = data["demand"]
22 capacity = data["capacity"]
23 vehicles = data["vehicles"]
24 edge_weight = data["edge_weight"]
25 service_time = data["service_time"]
26 time_window = data["time_window"]
27
28 model = Model("CVRP_PathLength")
29 model.setParam(’TimeLimit’, 50)
30
31 # Decision Variables
32 x = model.addVars(num_nodes, num_nodes, vtype=GRB.BINARY, name="x")
33 u = model.addVars(num_nodes, vtype=GRB.CONTINUOUS, name="u") # For capacity
34 arrival_time = model.addVars(num_nodes, vtype=GRB.CONTINUOUS, lb=0, name="arrival_time")
35
36 # Objective: Minimize total distance
37 model.setObjective(
38 quicksum(edge_weight[i][j] * x[i, j] for i in range(num_nodes) for j in range(num_nodes) if i != j),
39 GRB.MINIMIZE
40 )
41
42 # Core Constraints
43 model.addConstrs((quicksum(x[i, j] for i in range(num_nodes) if i != j) == 1 for j in range(1, num_nodes)),
44 "in_degree")
45 model.addConstrs((quicksum(x[i, j] for j in range(num_nodes) if i != j) == 1 for i in range(1, num_nodes)),
46 "out_degree")
47 model.addConstr(quicksum(x[0, j] for j in range(1, num_nodes)) <= vehicles, "max_vehicles")
48 model.addConstr(quicksum(x[0, j] for j in range(1, num_nodes)) == quicksum(x[j, 0] for j in range(1, num_nodes)),
49 "depot_flow")
50 # MTZ Capacity Constraints
51 model.addConstr(u[0] == 0, "u_depot")
52 for j in range(1, num_nodes):
[…]
55 for i in range(num_nodes):
56 if i != j:
57 model.addConstr(
58 u[j] >= u[i] + demand[j] - capacity * (1 - x[i, j]),
[…]
60 )
61
62 # Time window constraints
63 model.addConstr(arrival_time[0] == 0, name="StartTimeDepot")
64 model.addConstrs(
65 (arrival_time[i] + service_time[i] + edge_weight[i][j] <= arrival_time[j] + (1 - x[i, j]) * 1e6 for i in
66 range(num_nodes) for j in range(1, num_nodes) if i != j), name="TravelTime")
67 model.addConstrs((arrival_time[j] >= time_window[j][0] for j in range(1, num_nodes)), name="EarliestArrival")
68 model.addConstrs((arrival_time[j] <= time_window[j][1] for j in range(num_nodes)), name="LatestArrival")
69
70 # Solve
71 model.optimize()
72
73 # Core route extraction logic
74 routes = []
75 if hasattr(model, ’status’) and model.status in [GRB.OPTIMAL, GRB.TIME_LIMIT]:
76 visited = set()
77 for j in range(1, num_nodes): # Skip depot
78 if x[0, j].x > 0.5 and j not in visited:
79 current, route = j, []
80 while current != 0: # Until return to depot
81 route.append(current)
82 visited.add(current)
83 current = next((k for k in range(num_nodes) if k != current and x[current, k].x > 0.5), 0)
84 if route:
85 routes.append(route)
86 return routes
OR-tools Code:
1def find_feasible_routes(solution: VrpState) -> list[list[int]]:
2 """
3 Finds feasible routes for a VRP problem using an optimization solver.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "capacity": Maximum load capacity per vehicle.
9 - "demand": A 1D NumPy array of node demands (0 for the depot).
10 - "time_window": A 2D NumPy array where each row represents the [earliest start, latest end] time windows for servicing each node. The maximum time limit for each route is set to the depot’s latest end time.(0 for the depot)
11 - "service_time": A 1D NumPy array of node service times (0 for the depot).
12 - "edge_weight": A 2D NumPy array or matrix of distances between nodes.
13 Returns:
14 routes (list): A list of optimized routes, where each route is a list of node IDs (integers, excluding the depot node 0).
15
16 Note:
17 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
18 Ensure each node is visited exactly once.
19 """
20 # Data Preprocessing
21 data = solution.problem_data
22 # Ensure all data types are correct
23 edge_weight = data[’edge_weight’].astype(np.int64)
24 demands = [int(x) for x in data[’demand’].astype(np.int64)]
25 time_windows = [[int(tw[0]), int(tw[1])] for tw in data[’time_window’]]
26 service_times = [int(st) for st in data[’service_time’].astype(np.int64)]
27 vehicle_capacity = int(data[’capacity’])
28 num_nodes = len(data["demand"])
29 num_vehicles = num_nodes - 1
30
31 # Create the routing index manager
32 manager = pywrapcp.RoutingIndexManager(num_nodes, num_vehicles, 0)
33
34 # Create Routing Model
35 routing = pywrapcp.RoutingModel(manager)
36
37 # Create and register a transit callback for distance.
38 def distance_callback(from_index, to_index):
39 # Returns the distance between the two nodes.
40 from_node = manager.IndexToNode(from_index)
41 to_node = manager.IndexToNode(to_index)
42 return edge_weight[from_node][to_node]
43
44 transit_callback_index = routing.RegisterTransitCallback(distance_callback)
45
46 # Define cost of each arc.
47 routing.SetArcCostEvaluatorOfAllVehicles(transit_callback_index)
48
49 # Add Capacity constraint
50 def demand_callback(from_index):
51 # Returns the demand of the node.
52 from_node = manager.IndexToNode(from_index)
53 return demands[from_node]
54 demand_callback_index = routing.RegisterUnaryTransitCallback(demand_callback)
55 routing.AddDimensionWithVehicleCapacity(
56 demand_callback_index,
57 0, # null capacity slack
58 [vehicle_capacity] * num_vehicles, # vehicle maximum capacities
59 True, # start cumul to zero
60 ’Capacity’
61 )
62
63 # Add Time Window constraint
64 def time_callback(from_index, to_index):
65 # Returns the travel time between the two nodes (excluding service time).
66 from_node = manager.IndexToNode(from_index)
67 to_node = manager.IndexToNode(to_index)
68 return edge_weight[from_node][to_node] # Only travel time, no service time
69
70 time_callback_index = routing.RegisterTransitCallback(time_callback)
71 horizon = time_windows[0][1] # Maximum time limit for routes (depot’s latest end time)
72 routing.AddDimension(
73 time_callback_index,
74 horizon, # allow waiting time
75 horizon, # maximum time per vehicle
76 False, # don’t force start cumul to zero
77 ’Time’
78 )
79
80 # Set time windows for all nodes (including the depot)
81 time_dimension = routing.GetDimensionOrDie(’Time’)
82 for node_index in range(num_nodes): # Include depot node
83 index = manager.NodeToIndex(node_index)
84 time_dimension.CumulVar(index).SetRange(time_windows[node_index][0], time_windows[node_index][1])
85 # Add service time to the departure time
86 routing.solver().Add(
87 time_dimension.CumulVar(index) + service_times[node_index] <= time_dimension.CumulVar(index) + service_times[node_index]
88 )
89
90 # Solver Configuration
91 search_params = pywrapcp.DefaultRoutingSearchParameters()
92 search_params.first_solution_strategy = (
93 routing_enums_pb2.FirstSolutionStrategy.PATH_CHEAPEST_ARC
94 )
95 search_params.time_limit.seconds = 50 # Timeout protection
96
97 # Route Extraction
98 routes = []
99 if (result := routing.SolveWithParameters(search_params)):
100 for vehicle_id in range(routing.vehicles()):
101 index = routing.Start(vehicle_id)
102 route = []
103 while not routing.IsEnd(index):
104 node = manager.IndexToNode(index)
105 if node != 0: # Filter out depot node
106 route.append(node)
107 index = result.Value(routing.NextVar(index))
108 if route: # Filter out empty routes
109 routes.append(route)
110
111 return routes
Appendix G Database
The database contains six representative constraints, which exemplify problems related to these constraints. If no example problem is associated with the given input $I$, the base case $(I_{0},C_{0})$, which includes no additional constraints, will be selected. The following sections introduce these six representative constraints and the base case.
G.1 No Additional Constraints
Constraint Description: No additional constraints.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "demand": A 1D NumPy array of node demands (0 for the depot).
9 - "capacity": Maximum load capacity per vehicle.
10 - "edge_weight": A 2D NumPy array or matrix of distances between nodes.
11 - "service_time": A 1D NumPy array of node service times (0 for the depot).
12 - "time_window": A list of [earliest start, latest end] time windows for servicing each node.
13 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
14
15 Returns:
16 bool: True if the solution satisfies the constraints, False otherwise.
17
18 Note:
19 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
20 """
21 # Your code goes here...
22
23 return True
G.2 Vehicle Capacity
Constraint Description: The total load of each route must not exceed the vehicle capacity.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "edge_weight": A 2D NumPy array of distances between nodes.
9 - "demand": A 1D NumPy array of node demands (0 for the depot).
10 - "capacity": Maximum load capacity per vehicle.
11 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
12
13 Returns:
14 bool: True if the solution satisfies the constraints, False otherwise.
15
16 Note:
17 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
18 """
19 for route in solution.routes:
20 total_demand = sum(solution.problem_data["demand"][node] for node in route)
21 if total_demand > solution.problem_data["capacity"]:
22 return False
23 return True
G.3 Distance Limit
Constraint Description: Each route’s total length must not exceed 100 units.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "edge_weight": A 2D NumPy array of distances between nodes.
9 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
10
11 Returns:
12 bool: True if the solution satisfies the constraints, False otherwise.
13
14 Note:
15 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
16 """
17 max_route_length = 100
18 distances = solution.problem_data["edge_weight"]
19
20 for route in solution.routes:
21 tour = [0] + route + [0] # Add depot at the beginning and end of the route
22 route_length = sum(distances[tour[idx - 1]][tour[idx]] for idx in range(1, len(tour)))
23
24 # Check if the route length exceeds the maximum allowed length
25 if route_length > max_route_length:
26 return False
27
28 return True
G.4 Time Windows
Constraint Description: The arrival time at the node must meet the time windows.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "edge_weight": A 2D NumPy array of distances between nodes.
9 - "service_time": A 1D NumPy array of node service times (0 for the depot).
10 - "time_window": A list of [earliest start, latest end] time windows for servicing each node.
11 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
12
13 Returns:
14 bool: True if the solution satisfies the constraints, False otherwise.
15
16 Note:
17 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
18 """
19 for route in solution.routes:
20 current_time = 0 # Start at time 0
21 tour = [0] + route + [0] # Add depot at the beginning and end of the route
22
23 for idx in range(1, len(tour)):
24 arrive_time = current_time + solution.problem_data["edge_weight"][tour[idx - 1]][tour[idx]]
25 wait_time = max(0, solution.problem_data["time_window"][tour[idx]][0] - arrive_time)
26 current_time = arrive_time + wait_time
27
28 tw_start, tw_end = solution.problem_data["time_window"][tour[idx]]
29 if current_time > tw_end:
30 return False
31
32 # Add the service time for the current node after arriving and waiting
33 current_time += solution.problem_data["service_time"][tour[idx]]
34
35 return True
G.5 Pickup and Delivery
Constraint Description: At Node [24], 10 units of goods are picked up. Unlike delivery, visiting a pickup node reduces the vehicle’s available capacity.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "edge_weight": A 2D NumPy array of distances between nodes.
9 - "demand": A 1D NumPy array of node demands (0 for the depot).
10 - "capacity": Maximum load capacity per vehicle.
11 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
12
13 Returns:
14 bool: True if the solution satisfies the constraints, False otherwise.
15
16 Note:
17 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
18 """
19
20 pickup_nodes = {24: 10} # values in user requirements
21 vehicle_capacity = solution.problem_data["capacity"]
22
23 for route in solution.routes:
24 current_load = 0
25
26 for node in route:
27 demand = solution.problem_data["demand"][node]
28
29 if node in pickup_nodes:
30 current_load -= pickup_nodes[node] # When a pickup node is visited
31 else:
32 current_load += demand # If not a pickup point, it must satisfy the vehicle capacity.
33
34 # Check if the vehicle capacity is exceeded or negative after visiting this node
35 if current_load > vehicle_capacity or current_load < 0:
36 return False
37
38 return True
G.6 Same Vehicle
Constraint Description: Nodes [1, 10] must be served by the same vehicle.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "edge_weight": A 2D NumPy array of distances between nodes.
9 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
10
11 Returns:
12 bool: True if the solution satisfies the constraints, False otherwise.
13
14 Note:
15 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
16 """
17 for route in solution.routes:
18 if 1 in route and 10 in route:
19 break
20 else:
21 # If no route contains both nodes 1 and 10
22 return False
23
24 return True
G.7 Priority
Constraint Description: Nodes [17, 19] are priority points.
Verification Code:
1def check_constraints(solution: VrpState) -> bool:
2 """
3 Check if the solution satisfies the constraints.
4
5 Args:
6 solution (VrpState): An object representing the VRP solution.
7 - problem_data (dict): A dictionary with problem details:
8 - "edge_weight": A 2D NumPy array of distances between nodes.
9 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0).
10
11 Returns:
12 bool: True if the solution satisfies the constraints, False otherwise.
13
14 Note:
15 The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added.
16 """
17 # Define priority levels for nodes. Higher numbers mean higher priority.
18 priority_levels = {
19 17: 1, # Priority level 1 (highest priority)
20 19: 1, # Priority level 1 (highest priority)
21 }
22
23 # Default priority for non-listed nodes (lowest priority)
24 default_priority = 0
25
26 for route in solution.routes:
27 last_priority = max(priority_levels.values()) # Initialize to the highest priority
28
29 for node in route:
30 # Get current node’s priority, using the default if not specified
31 current_priority = priority_levels.get(node, default_priority)
32
33 # If we encounter a node with higher priority than the last, the constraint is violated
34 if current_priority > last_priority:
35 return False
36
37 # Update the last seen priority
38 last_priority = current_priority
39
40 return True